OpenStudy (anonymous):

A spinner has 9 equal sections labeled 1 through 9. What is the probability of spinning an odd number on the spinner and tossing heads on a coin?

OpenStudy (wolf1728):

Probability of odd number on spinner = 5/9 Probability of heads = 1/2

OpenStudy (anonymous):

what is the ans

OpenStudy (anonymous):

helpppp

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

probability of getting a odd number on spin wheel = \[\frac{ 5 }{ 9 }\] probability of getting head on the coin = \[\frac{ 1 }{ 2 }\] both are independent events. for independent events \[P(A \cap B) = P(A) \times P(B) = \frac{5 }{ 9 } \times \frac{ 1 }{ 2 } = \frac{ 5 }{ 18 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome :)